a coffee company wants a new flavor of cajan coffee. How many pounds of coffee worth $7 a pound should be added to 30 pounds of coffee worth $3 a pound to get a mixture worth $5 a pound
Answer by mananth(16946)

---------- dollars ---------------- quantity
Coffee A 3 ---------------- 30 pounds
Coffee B 7 ---------------- x pounds
Mixture 5 ---------------- 30 + x pounds

3 * 30 + 7 x = 5 ( 30 + x )
90 + 7 x = 150 + 5 x
7 x -5 x = -90 + 150
2 x = 60
/ 2
x = 30 pounds Coffee B

m.ananth@hotmail.ca